OpenStudy (anonymous):

medals to be given find the slope if a= (1,4) and b= (-2,3)

OpenStudy (freckles):

\[\text{ SLOPE going through points (x_1,y_1) and (x_2,y_2) }=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

so the answer is -1/3

OpenStudy (anonymous):

@frekles is that right

OpenStudy (anonymous):

@freckles is that correct

OpenStudy (freckles):

\[\frac{4-3}{1-(-2)}=\frac{1}{1+2}=\frac{1}{3}\] You might want to recheck because I got something else

OpenStudy (freckles):

\[\frac{3-4}{-2-1}=\frac{-1}{-3}=\frac{1}{3}\]

OpenStudy (freckles):

either way you did it you should gotten 1/3

OpenStudy (anonymous):

oh okay can you help me with